OpenStudy (anonymous):

if s= 7 cm and t=24cm what is the length of r?

OpenStudy (anonymous):

\[a^2 +b^2 = c^2\]

OpenStudy (anonymous):

its 25cm or 28cm or 31cm HEL[PPPPP IVE BEEN STUCK FOR 18 MINS ON THIS

OpenStudy (anonymous):

Apply Pythagoras' Theorem, which I posted above. \[7^2 + 24^2 = r^2\] What is 'r'?

OpenStudy (anonymous):

i think its28cm? @SmileyMan

OpenStudy (anonymous):

Show me your calculation steps.

OpenStudy (anonymous):

a2 + b2 = c2

OpenStudy (anonymous):

Yes, and?

OpenStudy (anonymous):

72+242=r2

OpenStudy (anonymous):

That is not correct. Remember: \[x^2=x*x \] So: \[7^2 = 7*7\]

OpenStudy (anonymous):

31!

OpenStudy (anonymous):

Show me your calculation steps.

OpenStudy (anonymous):

wait i dont get it

OpenStudy (anonymous):

wait no it is 25 for sure now!

OpenStudy (anonymous):

Pythagoras Theorem: \[a^2 + b^2 = c^2\]\[7^2+24^2=r^2\]\[7*7+24*24=r^2\]

OpenStudy (anonymous):

yup its 25 cm !